a)   What is the percentage change in the real price (1990 dollars) from 1980 to 2000? Compare this with your answer in (b). What do you notice? Explain.


1
Expert's answer
2021-10-26T08:55:49-0400

A significant part of the question is missing. However, this is how the answers would have appeared.



Reals price decrease "= 2.98 - 1.91 = 1.07"


The percentage change in real price from 1980 to 2000 is "=\frac{1.07}{2.98} * 100\% = 35.9\%"


The results are similar to those in part (b) despite the rounding-off error. The outcome shows that, regardless of the base year chosen when calculating the percentage changes in real price, the results have an insignificant variation.


Note the base year in this case "1990 = 100"